Let U be a universal set with disjoint subsets A and B; n(U) = 60, n(A) =25, and n(B) = 30. Find n((A U B)')
Answer by stanbon(75887)

Let U be a universal set with disjoint subsets A and B; n(U) = 60, n(A) =25, and n(B) = 30. Find n((A U B)')
-----
Since A and B are disjoint n(AUB) = n(A)+n(B) = 25+30 = 55
------
Suce n(U) = 60, n(AUB)' = 60-55 = 5
----------
Cheers,
Stan H.
-----------